# xTrimoPGLM: Unified 100B-Scale Pre-trained Transformer for Deciphering the Language of Protein

### v1

@@ 2.1.1 Zero-shot Naturalness
(a) Unified Objective to MLM
(b) Unified Objective to GLM
(c) MLM-adapted GLM vs. GLM (Scratch)
(d) GLM-adapted MLM vs. MLM (Scratch)
Figure 8: The empirical analysis of unified training.
Figure 9: Trials on different strategies for transition from Stage-1 to Stage-2.
Figure 10: More protein sequences visualizations generated by xTrimoPGLM-100B. (Left: Generated by xTrimoPGLM-100B. Right: Natural Proteins.)
Figure 11: The left panel illustrates the dataset composition used for pre-training the model. The right panel depicts the distribution of taxonomic categories of Uniref90, visualized as concentric circles representing the levels of superkingdom, kingdom, and phylum from innermost to outermost. The innermost circle represents four superkingdoms: Bacteria (67%), Archaea (3%), Eukarya (27%), and Viruses (1%), with 2% of sequences labeled as unclassified. The middle circle encompasses 17 classified kingdoms, including an unclassified bacteria category, denoted as “bacteria*”. The outermost circle denotes the phylum level, marking only those labels with counts over 200,000. In total, Uniref90 includes 273 known phyla.
Figure 12: Training Data Distribution
Figure 13: Scatter plots show delta perplexity for CAMEO and CASP15. The perplexities (PPL) are from the PLM modules of xT-Fold and ESMFold. Points represent proteins, with color gradients indicating perplexity delta by the x-axis PPL minus the y-axis PPL.
@@
Figure 15: Sequences comparison about generated antibodies.
@@
(a) Avg Seq. ID vs T
(b) Avg Seq. ID vs P
(c) Length vs T
(d) Length vs P
Figure 18: The sequence identity and length distributions of generated sequences with different temperatures ($T$) and nucleus sampling probability ($P$).
@@
 Downstream Task   Autoenc. Autoreg. Enc.-Dec. GLM
Example
 Protein Understanding
$\times$
Contact Prediction
 Protein Generation
$\times$
Antibody Re-design
@@
ESM-1b (?),
ESM2 (?),
Pro.BERT (?)
ProGen (?),
ProGen2 (?),
ProtGPT2 (?)
ProtTrans (?),
Ankh (?)
xTrimoPGLM
@@
Struc. Cont. P. Top L/5 ACC 12K 1.5K 1.5K ESM2-15B (?) ${}^{\clubsuit}$ 93.32
Fold. P. 12K-cls ACC 12.3K 736 3.2K Ankh_L (?) ${}^{\blacklozenge}$ 75.61
Sec. Struc. P. 3-cls ACC 11K - 39 Ankh_L (?) ${}^{\blacklozenge}$ 75.33
Dev. Sol. P. 2-cls ACC 62.4K - 6.9K ESM2-15B (?) ${}^{\clubsuit}$ 79.45
Stab. P. SRCC 53.6K 2.5K 12.8K ESM2-15B (?) ${}^{\clubsuit}$ 84.21
Temp. Stab. MCC 283K 63K 73.2K TemStaPro (?) ${}^{\blacklozenge}$ 94.22
Opt. Temp. SRCC 1.7K - 190 DeepET (?) 62.40 ${}^{\blacklozenge}$ 73.96
Opt. PH AUC 7.1K 760 1.9K ESM2-15B (?) 62.48 ${}^{\blacklozenge}$ 64.99
Clo. CLF AUC 23.3K - 4.7K ESM2-3B (?) 77.09 ${}^{\blacklozenge}$ 84.82
Mat. Pro. AUC 23.3K - 4.7K ESM2-15B (?) 79.17 ${}^{\blacklozenge}$ 86.48
Inter. Metal B. 2-cls ACC 6K - 1.3K ESM2-15B (?) ${}^{\clubsuit}$ 82.78
Enzyme Eff. PCC 13.5K 1.7K 1.7K DLKcat (?) ${}^{\blacklozenge}$ 74.79
Pept.-HLA Aff. AUC 575K 144K 171K CcBHLA (?) ${}^{\blacklozenge}$ 96.68
TCR-pMHC Aff. AUC 19.5K - 4.5K epiTCR (?) ${}^{\blacklozenge}$ 95.10
Func. Antib. Res. 19-cls ACC 2K - 1.3K ESM2-15B (?) ${}^{\clubsuit}$ 98.38
Fluor. P. SRCC 21.4K 5.4K 27.2K Ankh_L (?) 62.00 ${}^{\blacklozenge}$ 66.00
Fitness P. SRCC 6.3K 699 1.7K Ankh_L (?) 84.00 ${}^{\blacklozenge}$ 96.10
Loc. P. 10-cls ACC 6.6K - 1.8K Ankh_L (?) ${}^{\blacklozenge}$ 81.60
Table 3: Performance of different models across all benchmarked downstream protein-related tasks.
@@ H Chain L Chain Pair H Chain L Chain Pair
RMSDnormal-$\downarrow$ TM-scorenormal-$\uparrow$ DockQ normal-$\uparrow$ Inference time normal-$\downarrow$
@@
glu_activation
GeGLU
hidden dim.
10,240
ffn size
31,744
# layers
72
# attention heads
80
sequence_length
2,048
global batch size
4,224
max learning rate
4e-05
min learning rate
4e-06
adam_beta1
0.9
adam_beta2
0.95
adam_eps
1e-08
aggregated_samples_per_sequence
1,2,4,8
attention_dropout
0.1
attention_softmax_in_fp32
True
average_block_length
6
bias_dropout_fusion
True
checkpoint_activations
True
checkpoint_in_cpu
False
checkpoint_num_layers
9
clip_grad
1.0
tensor_parallel_size
4
pipeline_parallel_size
8
data_parallel_size
24
deepnorm
True
distributed_backend
nccl
eval_interval
300
fp16
True
mlm_prob
0.1
span_prob
0.2
gpt_prob
0.7
hidden_dropout
0.1
init_method_std
0.0052
initial_loss_scale
65536
layernorm_epsilon
1e-05
rotary_embedding
2D
learnable_rotary_embedding
False
length_per_sample
2048
log_interval
1
lr_decay_iter
None
lr_decay_samples
439,453,125
lr_decay_style
cosine
lr_warmup_samples
14,648,437
make_vocab_size_divisible_by
128
masked_softmax_fusion
True
micro_batch_size
1
min_gmask_ratio
0.4
min_loss_scale
1.0
optimizer
adamw
partition_activations
True
rampup_batch_size
240,24,12207031
save_interval
300
seed
1234
short_seq_prob
0.02
shrink_embedding_gradient_alpha
0.1
single_span_prob
0.02
split
949,50,1
tokenizer_type
ProteinTokenizer
weight_decay
0.1
zero_stage
1

### v2

@@ 8 Model FLOPs Comparisons
Figure 20: xTrimoPGLM generates a diverse set of sequences by varying temperature ($T\in{0.8,1.0,1.2,1.4,1.6}$) and nucleus sampling probability ($P\in{0.5,0.7,0.9,1.0}$) parameters. For each combination of $T$ and $P$, 600 sequences are sampled for comprehensive analysis. This figure displays the distributions of sequence identity (a, b) and length for the generated sequences across the different temperature and nucleus sampling probability settings. The box plots (c, d) mark the median, the orange horizontal line; upper and lower quartiles, purple horizontal line; upper and lower bound, black horizontal line; and 1.5$\times$ interquartile range (whiskers).
Table 1: The TM-score and GDT_TS score on CAMEO and CASP15.
Methods CAMEO (194) CASP15 (56)
TM-score GDT_TS TM-score GDT_TS
xT-Fold 0.86 0.83 0.70 0.64
ESMFold 0.85 0.81 0.65 0.59
OmegaFold 0.80 0.76 0.60 0.54
AF2 (Single Seq.) 0.39 0.28 0.33 0.23
AF2_model_3 (MSAs) 0.87 0.85 0.76 0.70
RoseTTAFold2 (MSAs) 0.86 0.83 0.73 0.68
@@
Downstream Task     Autoenc. Autoreg. Enc.-Dec. GLM Example
Protein Understanding     $\times$ Contact Prediction
Protein Generation     $\times$ Antibody Re-design
@@
ESM-1b (?), ESM2 (?), Pro.BERT (?) PROGEN (?), PROGEN2 (?), ProtGPT2 (?) ProtTrans (?), Ankh (?) xTrimoPGLM /
@@
Struc. Cont. P. Top L/5 ACC 12K 1.5K 1.5K ESM2-15B  (?) 92.19 93.32
Fold. P. 12K-cls ACC 12.3K 736 3.2K Ankh_L  (?) 61.10 75.61
Sec. Struc. P. 3-cls ACC 11K - 39 Ankh_L  (?) 80.70 75.33
Dev. Sol. P. 2-cls ACC 62.4K - 6.9K ESM2-15B  (?) 76.49 79.45
Stab. P. SRCC 53.6K 2.5K 12.8K ESM2-15B  (?) 80.75 84.21
Temp. Stab. MCC 283K 63K 73.2K TemStaPro  (?) 83.80 94.22
Opt. Temp. SRCC 1.7K - 190 DeepET  (?) 62.40 73.96
Opt. PH AUC 7.1K 760 1.9K ESM2-15B  (?) 62.48 64.99
Clo. CLF AUC 23.3K - 4.7K ESM2-3B  (?) 77.09 84.82
Mat. Pro. AUC 23.3K - 4.7K ESM2-15B  (?) 79.17 86.48
Inter. Metal B. 2-cls ACC 6K - 1.3K ESM2-15B  (?) 79.35 82.78
Pept.-HLA Aff. AUC 57.4K 7K 8.4K CcBHLA  (?) 95.00 96.68
TCR-pMHC Aff. AUC 19.5K - 4.5K epiTCR  (?) 92.50 95.10
Func. Antib. Res. 19-cls ACC 2K - 1.3K ESM2-15B  (?) 98.28 98.38
Fluor. P. SRCC 21.4K 5.4K 27.2K Ankh_L  (?) 62.00 66.00
Fitness P. SRCC 6.3K 699 1.7K Ankh_L  (?) 84.00 96.10
Loc. P. 10-cls ACC 6.6K - 1.8K Ankh_L  (?) 83.20 81.60
Enzyme Eff. PCC 13.5K 1.7K 1.7K DLKcat  (?) 71.00 74.79
Table 4: Performance of different models across all benchmarked downstream protein-related tasks. xT100B depicts xTrimoPGLM-100B model, E15B and E150M for ESM-15B and ESM-150M model respectively. Metric values are shown in both probing and LoRA (in parentheses) fine-tuning modes, where the underline denotes the best performance of probing and bold indicates the best performance of LoRA fine-tuning.
@@
Table 5: Performance comparisons between ProtT5  (?) with ESM2-15B, ESM2-3B, and xTrimoPGLM-100B on all eight downstream protein-related tasks via the Linear Probing approach. We utilize the best-performing protein language model, ProtT5-XL-U50, with 3 billion parameters. xT100B depicts xTrimoPGLM-100B model, E15B and E3B for ESM2-15B and ESM2-3B model respectively. The bold denotes the best performance and underline indicates the second-best performance.
Type Task Metric Model
xT100B E15B E3B ProtT5-XL3B
P. Struc. Cont. Pred. Top L/5 ACC 76.86 73.52 76.01 74.52
Fold Pred. 12-cls ACC 71.57 67.39 69.20 67.54
P. Func. Fluor. SRCC 65.16 63.84 56.68 52.67
Fitness SRCC 81.69 77.12 73.16 74.39
Localization 10-cls ACC 79.99 80.78 80.07 80.89
P. Inter. Metal Bind. 2-cls ACC 81.70 79.35 79.72 77.78
P. Dev. Solubility 2-cls ACC 76.04 74.76 72.74 74.91
Stability SRCC 75.52 71.69 69.93 73.26
Table 6: Comparison of xTrimo-series models across eight downstream protein-related tasks using the Linear Probing approach. We include models with 230M, 650M, 1B, 3B, 10B, and 100B parameters.Bold indicates the best performance, while underline signifies the second-best.
Type Task Metric Model
230M 650M 1B 3B 10B 100B
P. Struc. Cont. Pred. Top L/5 ACC 59.46 60.57 69.49 72.23 75.10 76.86
Fold Pred. 12-cls ACC 30.65 29.98 34.53 54.63 66.10 71.57
P. Func. Fluor. SRCC 54.59 45.59 55.62 56.65 58.85 65.00
Fitness SRCC 73.60 75.86 76.90 77.26 77.98 81.69
Localization 10-cls ACC 73.00 74.12 74.80 75.13 74.65 79.99
P. Inter. Metal Bind. 2-cls ACC 74.2 71.61 73.00 75.37 78.22 81.70
P. Dev. Solubility 2-cls ACC 71.21 72.63 73.43 74.00 74.56 76.04
Stability SRCC 74.20 73.70 74.61 80.10 78.36 75.52
Table 7: Comparison of performance (Spearman Coefficient) between zero-shot and supervised fitness prediction models. ’E1V’ refers to ESM-1V, while ’E650M’, ’E1B’, ’E3B’, and ’E15B’ denote ESM2 models with 650M, 1B, 3B, and 15B parameters, respectively. ’xT1B’, ’xT3B’, ’xT10B’, and ’xT100B’ represent xTrimo-series models. P.GEN2 denotes the PROGEN2 model. Results are based on single models, without ensemble methods. For ESM-1V, the average performance across its 5 model variants is reported.
Methods Zero-Shot (ProteinGym) Supervised (GB1)
- LP LoRA
ESM-1V 0.4045 - -
ESM2-650M 0.4172 0.7533 0.9515
ESM2-3B 0.4220 0.7669 0.9471
ESM2-15B 0.4169 0.7712 0.9475
P.GEN2-small 0.3223 - -
P.GEN2-base 0.3764 - -
P.GEN2-large 0.3761 - -
P.GEN2-xlarge 0.3781 - -
xT1B 0.4067 0.7690 0.9577
xT3B 0.4192 0.7726 0.9542
xT10B 0.4176 0.7798 0.9624
xT100B 0.3987 0.8169 0.9610
@@ H Chain L Chain Pair H Chain L Chain Pair
[…]
@@
glu_activation GeGLU
hidden dim. 10,240
ffn size 31,744
# layers 72
# attention heads 80
sequence_length 2,048
global batch size 4,224
max learning rate 4e-05
min learning rate 4e-06
adam_beta1 0.9
adam_beta2 0.95
adam_eps 1e-08
aggregated_samples_per_sequence 1,2,4,8
attention_dropout 0.1
attention_softmax_in_fp32 True
average_block_length 6
bias_dropout_fusion True
checkpoint_activations True
checkpoint_in_cpu False
checkpoint_num_layers 9
clip_grad 1.0
tensor_parallel_size 4
pipeline_parallel_size 8
data_parallel_size 24
deepnorm True
distributed_backend nccl
eval_interval 300
fp16 True
mlm_prob 0.1
span_prob 0.2
gpt_prob 0.7
hidden_dropout 0.1
init_method_std 0.0052
initial_loss_scale 65536
layernorm_epsilon 1e-05
rotary_embedding 2D
learnable_rotary_embedding False
length_per_sample 2048
log_interval 1
lr_decay_iter None
lr_decay_samples 439,453,125
lr_decay_style cosine
lr_warmup_samples 14,648,437
make_vocab_size_divisible_by 128
masked_softmax_fusion True
micro_batch_size 1
min_gmask_ratio 0.4
min_loss_scale 1.0
optimizer adamw
partition_activations True
rampup_batch_size 240,24,12207031
save_interval 300
seed 1234
short_seq_prob 0.02
shrink_embedding_gradient_alpha 0.1
single_span_prob 0.02
split 949,50,1
tokenizer_type ProteinTokenizer
weight_decay 0.1
zero_stage 1